6) 10 samples of 4140 steel (as received) were tensile tested from a population of 1000+ and the yield strengths had a frequency distribution as shown: Yield Strength (ksi) Frequency 72 69 65 64 58 What is the approximate mean yield strength of the population from which the samples were taken? a) 65.3 b) 65.6 c) 72.6 d) 82.0
Answer #1

Q.6) Let X = Yield Strength ( ksi ) and f = frequency

X f X * f
72 1 72 * 1= 72
69 3 69 * 3 = 207
65 2 65 * 2 = 130
64 2 64 * 2 = 128
58 2 58 * 2 = 116
f 10 x * f 653

Expected value of X is,

EX)24 *f 653 f10 653

The approximate mean yield strength of the population from which the samples were taken is 65.3

Answer: a) 65.3